In a right triangle, find the length of the side not given. Approximate to three decimal places, if necessary. Keep in mind that a and b are the lengths of the legs and c is the length of the hypotenuse.
b= 4 sqrt(2), c=9

Answer by nyc_function(2741)

Use a^2 + b^2 = c^2.
We are given b and c but not a. We are looking for side a.
a^2 + (4sqrt[2])^2 = 9^2
a^2 + 32 = 81
a^2 = 81 - 32
a^2 = 49
sqrt[a^2] = sqrt[49]
a = 7
The missing side is 7.